# OpenLLM-Ro - Technical Report on Open-source Romanian LLMs trained starting from Llama 2

### v1

@@ 1 Introduction
  • We present and release111huggingface.co/OpenLLM-Ro/RoLlama2-7b-Base-v1 the first foundational Romanian LLM (i.e., RoLlama) based on the open-source Llama 2 model Touvron et al. (2023). Our collection of models is grouped under the OpenLLM-Ro initiative.
  • We collect and translate a large set of conversation and instructions, resulting in RoLlama2-7b-Instruct222huggingface.co/OpenLLM-Ro/RoLlama2-7b-Instruct-v1 and RoLlama2-7b-Chat333huggingface.co/OpenLLM-Ro/RoLlama2-7b-Chat-v1 models capable of following instructions and entertaining multi-turn conversations in Romanian.
@@
  • We make our recipe (i.e., datasets and code) and models publicly available to encourage further research both for Romanian and other low or less resourced languages.
@@ 2 Related Work
With the advent of open source LLMs, especially Llama2 Touvron et al. (2023) and Mistral Jiang et al. (2023), non-English specialized alternatives have been trained.  Luukkonen et al. (2023) have trained a family of Finnish foundational models, ranging from 186M to 176B parameters. Trained on a total of 38B tokens (with a custom tokenizer), their models showcase state-of-the-art performance on a variety of Finish downstream tasks. Similarly, Cui et al. (2023) have trained both a foundational and a chat variant of Llama2 for Chinese, while Wang (2023) worked on a domain-specific (i.e., media) Chinese LLM. Finetuned chat-models starting from Llama2 or Mistral have been developed for Norwegian444huggingface.co/norallm, Dutch555huggingface.co/collections/BramVanroy/geitje-7b-ultra-65c1ee010ad80fd1f6a8f208, Bulgarian666huggingface.co/INSAIT-Institute, and Serbian777huggingface.co/gordicaleksa/YugoGPT.
@@ 3.2 Instructions and Conversations
We collect a large and diverse dataset of conversations and instructions to build the chat model. Since high-quality data in Romanian is very scarce, we resort to translating datasets such as Alpaca (both with GPT3.5888github.com/tatsu-lab/stanford_alpaca and GPT4999github.com/Instruction-Tuning-with-GPT-4/GPT-4-LLM), Dolly101010huggingface.co/datasets/databricks/databricks-dolly-15k, NoRobots111111huggingface.co/datasets/HuggingFaceH4/no_robots, GPT-Instruct121212github.com/teknium1/GPTeacher, Orca131313huggingface.co/datasets/Open-Orca/SlimOrca-Dedup, Camel141414huggingface.co/datasets/camel-ai/math151515huggingface.co/datasets/camel-ai/chemistry161616huggingface.co/datasets/camel-ai/biology171717huggingface.co/datasets/camel-ai/physics, UltraChat181818huggingface.co/datasets/stingning/ultrachat/tree/main, and OpenAssistant Köpf et al. (2023)191919huggingface.co/datasets/OpenAssistant/oasst1. Additionally, we manually translated the MTBench corpora Zheng et al. (2023), leading to a total of around 2.3M instructions and conversations that were translated202020www.systransoft.com/ last accessed 6th of March 2024 and used for our experiments.
@@ 4 Training
Training the foundational and chat variants is done using the official Llama repository212121github.com/facebookresearch/llama-recipes. We pick a fixed learning rate from early experiments with a small portion of the data instead of the cosine scheduler proposed in the original paper. Due to hardware limitations, we experiment with a maximum sequence length of 512, 1024, and 2048, picking the biggest batch size that fits into the memory (i.e., 64 for the sequence length of 512, 32 for the sequence length of 1024, and 16 for the sequence length of 2048). We use the proposed AdamW optimizer, with weight decay set to 0.1 and gradient clipping. We do not employ LORA Hu et al. (2021) in any stage, as we perform full training for both the continual pre-training and supervised finetuning steps.
@@ 4.1 Foundational Model
Training the foundational model is done iteratively using continual pretraining on the CulturaX dataset. Starting from the base Llama2222222huggingface.co/meta-llama/Llama-2-7b-hf model, we train the first model on 5% of the CulturaX, the following model on the next 5% (leading to the second model seeing 10% of the data), and the third model on another 10% of the CulturaX. Therefore, we have three variants of a Romanian foundational model, variants that used 5%, 10%, and 20% of the entire CulturaX dataset. Training the foundational model is done over one epoch, with a fixed learning rate of 1e-4.
@@ 5 Evaluation
Evaluating a general LLM is not a straightforward task considering the vast number of benchmarks and evaluation protocols proposed to judge the quality of a model. Popular benchmarks include Open LLM Leaderboard232323huggingface.co/spaces/HuggingFaceH4/open_llm_leaderboard, an online leaderboard that encompasses six reasoning and general knowledge tasks: ARC Clark et al. (2018), HellaSwag Zellers et al. (2019), MMLU Hendrycks et al. (2020), TruthfulQA Lin et al. (2021), Winogrande Sakaguchi et al. (2021) and GSM8k Cobbe et al. (2021). MT-Bench Zheng et al. (2024) proposes a multi-turn question set and shows that strong LLMs (e.g., GPT-4) can act as judges and can approximate human preferences in a scalable and somewhat explainable way. AGIEval Zhong et al. (2023) is a benchmark designed to evaluate foundational models in the context of human-centric standardized exams. Evaluation is done in multiple settings, from zero and few-shot to Chain of Thought (COT) prompting. While benchmarks in the Open LLM Leaderboard select the most probable answer (out of four candidates) for each question, the focus on MT-Bench is more on the quality of the generated text than the intrinsic knowledge of the world of each model. Some experiments242424twitter.com/maximelabonne/status/1746317053143773628?s=19 try and compute a correlation factor between human preferences and results of previously mentioned benchmarks.
Evaluating LLMs in any language besides English only adds another layer of complexity. For Romanian versions of ARC, HellaSwag, MMLU, and TruthfulQA, we resort to ChatGPT translations Dac Lai et al. (2023), while we translate252525same system used for translating instructions Winogrande and GSM8k. The generally used few-shot values for ARC, HellaSwag, and MMLU (i.e., 25 shots for ARC, 10 shots for HellaSwag, and 5 shots for MMLU) lead to the prompt having a length equal to or even surpassing 2048 tokens. As some of our models have been trained only on sequences of length 512, 1024, or 2048, we opt to use multiple few-shot settings. More specifically, for ARC, we test with 0, 1, 3, 5, 10, and 25 shots, MMLU with 0, 1, 3, and 5, HellaSwag with 0, 1, 3, 5, and 10 shots, and Winogrande is evaluated in 0, 1, 3 and 5 shot setting. For GSM8k, we employ 1, 3, and 5 shot evaluation. Results in Tables 1, 2 are averaged over a few shot settings for each benchmark. More detailed results are presented in the Appendix.
@@ 6 Results
We observe a very high performance of Mistral models (both foundational and chat). However, since the authors of Mistral (Jiang et al. (2023)) do not provide many details about their training data, we cannot know if the model has been trained on Romanian data as well as other languages. Furthermore, it seems that finetuning Mistral on Romanian Alpaca 262626huggingface.co/Andrei481/Mistral-7B-v0.1-Romanian/discussions/1 drastically decreases the performance on considered benchmarks.
Turning to an evaluation based more on generation rather than just picking the most probable continuation, the performance on MT-Bench is presented in Table 3. We note that even when using a Romanian prompt, most Llama2 and Mistral completions or answers are either entirely in English or have the first few words in Romanian, then reverting to English. Specifically asking the models to answer in Romanian in the prompt somewhat alleviates this issue, especially for Mistral models. As seen in the last column in Table 3, when specifically prompted to answer in Romanian, Llama2 manages to generate text in Romanian in only 25% of cases, while Mistral answers are almost entirely in Romanian (over 90% of cases vs 70% when not prompted to answer in Romanian). We automatically detect the language of the answer [^27] https://pypi.org/project/langdetect/ and overwrite with 0 the score of each answers that is not in Romanian.
@@ 7 Conclusions
Furthermore, we present a general recipe which we expect to work on different, more powerful LLMs as well (e.g., bigger Llama2, Mistral, Mixtral [^28] https://mistral.ai/news/mixtral-of-experts/ or Llama3292929https://ai.meta.com/blog/meta-llama-3/ models). We firmly believe that this is just the first step towards building and adapting LLMs for Romanian for both research and industry applications.
@@ Acknowledgements
The universities contributed with researchers who worked pro-bono (POLITEHNICA Bucharest and the University of Bucharest) and computing power for training the models (POLITEHNICA Bucharest). Synchronization and collaboration took place within a project [^30] https://ilds.ro/llm-for-romanian carried out in the Institute of Logic and Data Science with funding from BRD Groupe Societe Generale.

### v2

@@ 1 Introduction
  • We present and release111www.huggingface.co/OpenLLM-Ro/RoLlama2-7b-Base-v1 the first foundational Romanian LLM (i.e., RoLlama) based on the open-source Llama 2 model Touvron et al. (2023). Our collection of models is grouped under the OpenLLM-Ro initiative.
  • We collect and translate a large set of conversation and instructions, resulting in RoLlama2-7b-Instruct222www.huggingface.co/OpenLLM-Ro/RoLlama2-7b-Instruct-v1 and RoLlama2-7b-Chat333www.huggingface.co/OpenLLM-Ro/RoLlama2-7b-Chat-v1 models capable of following instructions and entertaining multi-turn conversations in Romanian.
@@
  • We make our recipe444www.github.com/OpenLLM-Ro (i.e., datasets and code) and models publicly available to encourage further research both for Romanian and other low or less resourced languages.
@@ 2 Related Work
With the advent of open source LLMs, especially Llama2 Touvron et al. (2023) and Mistral Jiang et al. (2023), non-English specialized alternatives have been trained.  Luukkonen et al. (2023) have trained a family of Finnish foundational models, ranging from 186M to 176B parameters. Trained on a total of 38B tokens (with a custom tokenizer), their models showcase state-of-the-art performance on a variety of Finish downstream tasks. Similarly, Cui et al. (2023) have trained both a foundational and a chat variant of Llama2 for Chinese, while Wang (2023) worked on a domain-specific (i.e., media) Chinese LLM. Finetuned chat-models starting from Llama2 or Mistral have been developed for Norwegian555www.huggingface.co/norallm, Dutch666www.huggingface.co/collections/BramVanroy/geitje-7b-ultra-65c1ee010ad80fd1f6a8f208, Bulgarian777www.huggingface.co/INSAIT-Institute, and Serbian888www.huggingface.co/gordicaleksa/YugoGPT.
@@ 3.2 Instructions and Conversations
We collect a large and diverse dataset of conversations and instructions to build the chat model. Since high-quality data in Romanian is very scarce, we resort to translating datasets such as Alpaca (both with GPT3.5999www.github.com/tatsu-lab/stanford_alpaca and GPT4101010www.github.com/Instruction-Tuning-with-GPT-4/GPT-4-LLM), Dolly111111www.huggingface.co/datasets/databricks/databricks-dolly-15k, NoRobots121212www.huggingface.co/datasets/HuggingFaceH4/no_robots, GPT-Instruct131313www.github.com/teknium1/GPTeacher, Orca141414www.huggingface.co/datasets/Open-Orca/SlimOrca-Dedup, Camel151515www.huggingface.co/datasets/camel-ai/math161616www.huggingface.co/datasets/camel-ai/chemistry171717www.huggingface.co/datasets/camel-ai/biology181818www.huggingface.co/datasets/camel-ai/physics, UltraChat191919www.huggingface.co/datasets/stingning/ultrachat/tree/main, and OpenAssistant Köpf et al. (2023)202020www.huggingface.co/datasets/OpenAssistant/oasst1. Additionally, we manually translated the MTBench corpora Zheng et al. (2023), leading to a total of around 2.3M instructions and conversations that were translated212121www.systransoft.com/ last accessed 6th of March 2024 and used for our experiments.
@@ 4 Training
Training the foundational and chat variants is done using the official Llama repository222222www.github.com/facebookresearch/llama-recipes. We pick a fixed learning rate from early experiments with a small portion of the data instead of the cosine scheduler proposed in the original paper. Due to hardware limitations, we experiment with a maximum sequence length of 512, 1024, and 2048, picking the biggest batch size that fits into the memory (i.e., 64 for the sequence length of 512, 32 for the sequence length of 1024, and 16 for the sequence length of 2048). We use the proposed AdamW optimizer, with weight decay set to 0.1 and gradient clipping. We do not employ LORA Hu et al. (2021) in any stage, as we perform full training for both the continual pre-training and supervised finetuning steps.
@@ 4.1 Foundational Model
Training the foundational model is done iteratively using continual pretraining on the CulturaX dataset. Starting from the base Llama2232323www.huggingface.co/meta-llama/Llama-2-7b-hf model, we train the first model on 5% of the CulturaX, the following model on the next 5% (leading to the second model seeing 10% of the data), and the third model on another 10% of the CulturaX. Therefore, we have three variants of a Romanian foundational model, variants that used 5%, 10%, and 20% of the entire CulturaX dataset. Training the foundational model is done over one epoch, with a fixed learning rate of 1e-4.
@@ 5 Evaluation
Evaluating a general LLM is not a straightforward task considering the vast number of benchmarks and evaluation protocols proposed to judge the quality of a model. Popular benchmarks include Open LLM Leaderboard242424www.huggingface.co/spaces/HuggingFaceH4/open_llm_leaderboard, an online leaderboard that encompasses six reasoning and general knowledge tasks: ARC Clark et al. (2018), HellaSwag Zellers et al. (2019), MMLU Hendrycks et al. (2020), TruthfulQA Lin et al. (2021), Winogrande Sakaguchi et al. (2021) and GSM8k Cobbe et al. (2021). MT-Bench Zheng et al. (2024) proposes a multi-turn question set and shows that strong LLMs (e.g., GPT-4) can act as judges and can approximate human preferences in a scalable and somewhat explainable way. AGIEval Zhong et al. (2023) is a benchmark designed to evaluate foundational models in the context of human-centric standardized exams. Evaluation is done in multiple settings, from zero and few-shot to Chain of Thought (COT) prompting. While benchmarks in the Open LLM Leaderboard select the most probable answer (out of four candidates) for each question, the focus on MT-Bench is more on the quality of the generated text than the intrinsic knowledge of the world of each model. Some experiments252525www.twitter.com/maximelabonne/status/1746317053143773628?s=19 try and compute a correlation factor between human preferences and results of previously mentioned benchmarks.
Evaluating LLMs in any language besides English only adds another layer of complexity. For Romanian versions of ARC, HellaSwag, MMLU, and TruthfulQA, we resort to ChatGPT translations Dac Lai et al. (2023), while we translate262626same system used for translating instructions Winogrande and GSM8k. The generally used few-shot values for ARC, HellaSwag, and MMLU (i.e., 25 shots for ARC, 10 shots for HellaSwag, and 5 shots for MMLU) lead to the prompt having a length equal to or even surpassing 2048 tokens. As some of our models have been trained only on sequences of length 512, 1024, or 2048, we opt to use multiple few-shot settings. More specifically, for ARC, we test with 0, 1, 3, 5, 10, and 25 shots, MMLU with 0, 1, 3, and 5, HellaSwag with 0, 1, 3, 5, and 10 shots, and Winogrande is evaluated in 0, 1, 3 and 5 shot setting. For GSM8k, we employ 1, 3, and 5 shot evaluation. Results in Tables 1, 2 are averaged over a few shot settings for each benchmark. More detailed results are presented in the Appendix.
@@ 6 Results
We observe a very high performance of Mistral models (both foundational and chat). However, since the authors of Mistral (Jiang et al. (2023)) do not provide many details about their training data, we cannot know if the model has been trained on Romanian data as well as other languages. Furthermore, it seems that finetuning Mistral on Romanian Alpaca 272727www.huggingface.co/Andrei481/Mistral-7B-v0.1-Romanian/discussions/1 drastically decreases the performance on considered benchmarks.
Turning to an evaluation based more on generation rather than just picking the most probable continuation, the performance on MT-Bench is presented in Table 3. We note that even when using a Romanian prompt, most Llama2 and Mistral completions or answers are either entirely in English or have the first few words in Romanian, then reverting to English. Specifically asking the models to answer in Romanian in the prompt somewhat alleviates this issue, especially for Mistral models. As seen in the last column in Table 3, when specifically prompted to answer in Romanian, Llama2 manages to generate text in Romanian in only 25% of cases, while Mistral answers are almost entirely in Romanian (over 90% of cases vs 70% when not prompted to answer in Romanian). We automatically detect the language of the answer282828www.pypi.org/project/langdetect/ and overwrite with 0 the score of each answers that is not in Romanian.
@@ 7 Conclusions
Furthermore, we present a general recipe which we expect to work on different, more powerful LLMs as well (e.g., bigger Llama2, Mistral, Mixtral292929www.mistral.ai/news/mixtral-of-experts/ or Llama3303030ai.meta.com/blog/meta-llama-3/ models). We firmly believe that this is just the first step towards building and adapting LLMs for Romanian for both research and industry applications.
@@ Acknowledgements
The universities contributed with researchers who worked pro-bono (POLITEHNICA Bucharest and the University of Bucharest) and computing power for training the models (POLITEHNICA Bucharest). Synchronization and collaboration took place within a project313131www.ilds.ro/llm-for-romanian carried out in the Institute of Logic and Data Science with funding from BRD Groupe Societe Generale.